Meena is thrice as old as Sheraz. Fatima will be twice as old as Meena 6 years hence. Six years ago Sheraz was five year old.What is Fatima's present age?
Answer by stanbon(75887)

Meena is thrice as old as Sheraz. Fatima will be twice as old as Meena 6 years hence. Six years ago Sheraz was five year old.What is Fatima's present age?
--------------
m = 3s
f+6 = 2(m+6)
s-6 = 5
---------
s = 11
m = 3*11 = 33
----
f = 2m+12-6
f = 2m+6
Fatima = 2*33+6 = 72 yrs.
---------------
Cheers,
Stan H.
----------------